# Challenges and Opportunities for Visual Analytics in Jurisprudence

### v2

@@ 6.3 Analytical Reasoning
The traditional UI of legal databases merely lists groups of legal documents related to the currently displayed document. There are no means by which legal scholars can add evidence of their insights to that list. In VA, analytical reasoning is a knowledge-generation process that can have systematic aspects but is also serendipitous [58]. Users must be aware of the evidence found throughout this process to increase the chance of insights. A prominent approach to capture such evidence is analytical provenance as suggested by Pirolli and Card for their sense-making loop [93]. For that, Pérez-Messina et al. suggest organizing analytical artifacts like annotations and their relationships [58, 113].
In related, text-based research practice, Overview allows investigative journalists to create tags that attach their findings to clusters [114]. Tian et al. add to this, allowing users to adapt clusters of literature interactively, adding comments about their understanding of topics [60] (refer to Figure 2, c). With ForceSPIRE, Endert et al. address visual text analytics more broadly, enabling experts to create text highlights that constitute a spatial layout of document relationships [107] (refer to Figure 2, b).
Conveying the relationships within and between such clusters is central for analytical reasoning, so node-link diagrams are frequently employed as visualization methods. Lettieri et al. introduce the Norm Graph Navigator, allowing interactive exploration of legal relationships [74]. More recently, La Cava et al. integrate natural language querying into network navigation, enabling users to filter nodes efficiently [75] (refer to Figure 2, c). Since analytical artifacts externalize tacit knowledge as explicit [48], they can act as a countermeasure to the confirmation bias in hypothesis generation [93] and enable collaboration [113], which can reinforce diverse opinions in legal decision-making. When persisted [113], these artifacts can help users recall and replicate insights [115]. Analytical provenance additionally fosters transparency, which is particularly important in the legal domain, where the validity and reliability of insights must be meticulously documented. Integrating these aspects, for example, through 2D clustering for overview, graph networks for relations, and provenance trees for documentation, shows the significant potential of VA for jurisprudence.
@@ 7 Emergent Challenges for Visual Analytics Design
While numerous approaches exist for data navigation in the legal domain, knowledge representation and analytical reasoning are often left to analysts to perform manually, as we have seen in the previous discussion. We have analyzed several examples of existing VA techniques to address these shortcomings in section 6. Crucially, no existing solution integrates all these areas of VA into a unified framework to exploit their synergetic effects. In the following, we explore the opportunities and challenges of developing such an integrated approach, building upon the proposed solutions.
@@ 7.1 Addressing Jurisprudence
Traditionally, German legal scholars predominantly use keyword-based search interfaces, the prevalent UI paradigm [26, 27]. However, this poses several challenges to legal scholars, hindering effective workflow support. While familiar, the paradigm limits scholars to primitive input, lacking the expressiveness required in the legal domain. Hence, we propose combining existing visualization techniques into a novel VA workflow (refer to Figure 1) for jurisprudence, leveraging LLM-assisted contextual understanding.
@@
As part of the initial workflow phase (refer to $W_{\texttt{P1}}$), the scholar scopes their research verbally. Instead of a syntactic keyword search, they articulate analysis goals using natural language. The system processes the verbalization and embeds the query into a semantic context by inferring references to legal entities such as laws, legal norms, and court rulings. We imagine an overview displaying clusters of legal concepts and individual documents relevant to the query, enabling scholars to leverage their domain expertise to refine the selection. For its implementation, the overview leverages a network visualization augmented with colored shapes that indicate clusters. For details, a scholar can preview individual documents on-demand during selection, highlighting explanations for their relevance. After refining the scoping, legal scholars transition to the second workflow phase (refer to $W_{\texttt{P1}}$). In VA, the UI transitions into a detailed view upon selecting a cluster or document of interest. We propose juxtaposing the formal legal structure with a branching visualization such as an Icicle Plot. The former displays the legal documents of the scope with their hierarchy, enabling efficient navigation. Meanwhile, the Icicle Plot provides the contextualized path of the selected document in its hierarchy. Together, these visualizations address the legal domain’s unique requirement for simultaneous navigation across multiple hierarchies.
@@
During the analysis & interpretation, legal reasoning takes place, which must utilize the scholar’s domain knowledge. While the proposed VA techniques can initially visualize the explicit knowledge from the legal documents, they cannot draw on evolving domain knowledge. To achieve this, VA must provide interaction for users to articulate their tacit knowledge. Therefore, at any point during the workflow, if users find a relevant document missing from the current selection, they can establish a relationship to include it. The Treemap and Icicle Plot visualizations facilitate this interaction by integrating empty cells with suggestive icons as visual placeholders. By clicking these icons, the analyst can relate a selected document to others represented in the visualizations. These interactions reveal user preferences relevant to the subsequent recommendation of related legal documents. With these user preferences, we augment a knowledge graph (KG) of the legal documents selected during the first workflow phase. Yang et al. demonstrate the benefit of such an augmentation with LegalGNN, a framework for recommending legal documents based on a graph neural network trained with a legal KG and the contents of legal documents [116]. We exploit the resulting recommendations to personalize the VA workflow with the expert’s domain knowledge. Further, we draw tacit knowledge from its externalization through analytical artifacts.
@@
Besides the Treemap and the Icicle Plot, we propose to add a concept map to the top side of the suggested user interface. The map allows users to synthesize and document evidence acquired throughout the legal reasoning process, forming a crucial part of the third workflow phase (refer to $W_{\texttt{P3}}$). Scholars capture and connect relevant artifacts within the concept map to organize their legal insights. Nodes represent primitive artifacts resembling individual legal entities, such as legal norms or court rulings. In contrast, compound artifacts describe the links that relate multiple primitives, representing causal connections in the legal realm. Based on these and the graph augmented during the previous workflow phase, we enhance documentation of the legal reasoning process through retrieval-augmentation generation (RAG). Edge et al. propose GraphRAG, which performs RAG on the KG of text corpora [117]. The authors leverage the hierarchy of the KG to enable the summarization of documents at varying levels of abstraction. This approach aligns with the nature of legal corpora and can benefit the documentation of legal reasoning in jurisprudence. For instance, in analyzing racial profiling by the German Federal Police (refer to subsection 5.1), a scholar might document that the police are responsible for securing train stations (§ 3 BPolG) and are permitted to identify individuals if necessary for fulfilling their duties (§ 23 BPolG). However, this authority must be balanced against the third-party effect of the German Basic Law (Art. 3 GG), which ensures equality and non-discrimination.
@@
Implementing the proposed VA workflow faces challenges. The suggested visualization techniques for VA may suffer from issues with visual scalability. For vast legal datasets, node-link diagrams, Treemaps, and Icicle Plots are of limited use. However, since the datasets are hierarchically structured, we can progressively unveil parts of the hierarchy to limit the information these visualization techniques need to convey. Further, leveraging progress in natural language understanding remains difficult: legal text data combines domain-specific language [4, p. 123] and precise phraseology with intentional ambiguity at times [4, p. 128]. This ambiguity allows for statutory interpretation [4, p. 22], enabling laws to evolve alongside societal norms and values, but is challenging for LLMs. Although trained on vast datasets, Large Language Models still struggle to apply these nuances of legal reasoning, particularly open foundation models. For example, we found that Llama 3.1 70B still sometimes confuses § 212 StGB (Strafgesetzbuch, i.e., criminal code), which defines homicide for murder (§ 211 StGB). However, by definition of the criminal code, they are two strictly separate criminal offenses. While such models can describe the differences in great detail when explicitly asked, during reasoning, they can tend to collapse any distinctions without further warning. Equally, hallucinations, particularly during the subsumption phase, are critical as LLMs struggle with completeness. Therefore, it is crucial to perform extensive fine-tuning of foundational LLMs to align them with the specific language of the legal domain. This also presents ample opportunity for VA design to facilitate collaboration between human scholars and LLMs during the legal reasoning process.
@@ 7.2 Implications for Visual Analytics Design
The lingering challenges in data navigation, knowledge externalization, and analytical reasoning cannot be sufficiently addressed with AI and computational solutions alone. Therefore, we advocate for a human-is-the-loop design [118], which positions the VA system as an integral part of the scholar’s workflow. This view contrasts designs that rely on the common human-in-the-loop [15, 47] approach. While the latter assumes the domain expert to become a part of their VA system, the former prerequisites the VA system to become part of the domain expert and their workflow instead. To achieve this, the VA design process requires a user-centered approach that honors the workflows of the domain expert, supporting them throughout the sense-making loop [93, 118]. Therefore, we highlight the following key techniques and higher-level considerations regarding jurisprudence and generalization to other domains.
@@
User-Centered Design — The development of effective VA design requires adherence to user-centered design principles [52, 57]. These principles promote broader adoption of VA designs that meet specific audience needs. These needs vary across different application domains, valuing various visual qualities, such as page layout in law. While these aspects may not seem directly related to the effectiveness or efficiency of VA design, they can significantly influence data navigation in information spaces. This variation in user requirements presents ample opportunity for employing design studies or conducting pilot interviews with domain experts. These methodological approaches allow for gaining critical insights into the audience’s specific needs, workflows, and mental models, like the importance of tacit knowledge, the lack of contextualized search, and the three-phase research process, which we discussed in section 4.
@@
Knowledge Externalization — In application domains, subject matter experts possess tacit knowledge regarding domain-specific problems. It encompasses implicit understanding, experiential insights, and intuitive reasoning processes not readily encoded in explicit knowledge. The externalization process captures these domain-specific analytical approaches in shareable analytical artifacts. If not articulated, VA design inherently lacks access to this knowledge, potentially compromising analytical effectiveness. While explicit knowledge may be available in substantial quantities through text documents, for example, integrating with tacit knowledge is imperative to enhance analytical outcomes [45]. VA design must strive to externalize tacit knowledge through user interaction, KGs, and analytical artifacts.
@@
Provenance Tracking — Analytical artifacts manifest knowledge externalization processes within VA systems. These artifacts serve as persistent representations of analytical reasoning and decision-making trajectories throughout the analytical workflow [113]. Since users often struggle to maintain holistic awareness of past analytical steps, access to provenance information is essential for branching and revising the analysis process. VA design must facilitate the creation and storage of analytical artifacts through comprehensive provenance tracking, fostering the user’s awareness of the analysis process. Concurrently, these artifacts can be systematically utilized to guide subsequent analytical processes. VA design must exploit this recursive employment of analytical artifacts to create a self-reinforcing system that progressively refines the analytical workflow based on accumulated insights and knowledge.
@@
Natural Language Inference — While mining user interactions with VA interfaces can constitute guidance, it presents inherent methodological challenges in deducing higher-order analytical objectives [59]. Interaction patterns alone provide insufficient semantic context for accurately inferring analytical goals that guide user behavior. With recent advancements in LLMs, these models demonstrate sophisticated semantic parsing and intention recognition capabilities that can be leveraged to extract structured analytical objectives from natural language [119]. VA design must enable users to articulate these analytical objectives through natural language interfaces to capture nuanced analytical intentions while going beyond a simple chat-based interface design.
@@
Unified Design — To achieve effectiveness and efficiency in addressing domain-specific problems, VA design must integrate the aforementioned key techniques. User-centered design facilitates identifying and formalizing domain-specific workflows, as demonstrated through our structured interviews with subject matter experts (refer to section 4). Such approaches reveal critical insights into domain-specific analytical processes that may not be apparent through conventional design approaches. Likewise, knowledge externalization provides systematic access to tacit knowledge inherent to analysts. This externalization process transforms implicit knowledge into explicit, computationally accessible representations that can inform VA. Provenance tracking can effectively reflect domain-specific workflows by capturing the structure of analytical processes. Minimizing cognitive load induced by context switching between applications resulting from fragmented VA design is imperative. Instead, it must unify analytical functionalities within a coherent interface to optimize cognitive resource utilization and enhance analytical performance.
@@ 7.3 Limitations
In this unified design, we focus on legal tech from the perspective of VA while adopting a definition of visualization [120, p. 1] from the visualization community in our literature analysis. Meanwhile, the legal community borrows a broader definition of visualization [65]. This difference leads to papers from the legal community that visualize legal data but are not considered VA. Although these approaches are not the primary focus of our work, we incorporate their visualization concepts, which are familiar to legal scholars, in our proposed VA design (refer to section 6). For example, we use Concept Maps, which are commonly used in the legal community to visualize cases and causal relationships between legal norms (e.g., Cantatore and Stevens [66]), to capture tacit knowledge (refer to subsection 6.2) during legal reasoning. Thus, visualization concepts from the legal domain can be incorporated into VA design, fostering interdisciplinary research. To enhance our understanding of how legal scholars operate, we describe different levels of tasks (refer to subsection 3.2). These levels are supported by research on applications in legal tech, often emerging from the intersection of law and computer science. Consequently, the labels for these task levels resemble technical terms from VA, such as information retrieval. However, these tasks are grounded in the operations of scholars as described in literature from the legal domain (refer to subsection 3.1). Additionally, these tasks are linked to the three-phase workflow (refer to subsection 5.1), which is based on interviews with legal scholars (refer to section 4).
@@
These results of the subject matter expert interviews (refer to section 4) take common workflows and recurring challenges faced by German scholars from various legal disciplines into account. Still, they are not exhaustive, as further disciplines exist in the German legal system and beyond. To accommodate the differences between legal disciplines, we have decided on a semi-structured interview, where we adapt the follow-up questions to the interviewees’ backgrounds and experiences. Although we assume differences in the details, we expect the significant challenges distilled in section 6 to be similar as they concern all disciplines. We elicit our observations from subject matter experts who work exclusively with the German civil law system and European legislation. So far, we have not incorporated experiences from countries that practice civil law or other legal systems like common law (refer to section 1), which is practiced in the United States, for example.
To illustrate the three-phase workflow design (refer to subsection 5.1) and to critically discuss traditional UI for legal research in our case study (refer to section 6), we have focused on the Bundespolizeigesetz (Federal Police Act). This specific legal research context highlights the challenges described in a broader context by participants during the interviews (refer to section 4) using a concrete example. Challenges arise from accessibility issues, navigational complexity, the inefficiency of the search interface, and the reliance on tacit knowledge, which extend to other laws and legal research contexts. This generalization allows our case study to transcend the specific example chosen and the German legal system. To illustrate the traditional UI used in jurisprudence as part of our case study (refer to section 6), we use beck-online as an example. Although beck-online is just one of many commercial and scientific applications for jurisprudence, it is a suitable representative. In the German legal system context, beck-online is among the most widely used legal research applications by practitioners. Additionally, beck-online features a design similar to widely used applications such as Westlaw [28] and PACER [121], which helps generalize our findings to case law. While we have theoretically examined the workflow design in subject matter expert interviews (refer to section 4) for various legal disciplines, our work lacks empirical evaluations of the proposed VA workflow in these specific settings. We partially rely on empirical evaluations of visualization techniques that we propose for use in a legal setting. However, the unified design incorporating these techniques has not been evaluated accordingly. Furthermore, the existing empirical evaluations have been conducted in related domains but may not necessarily generalize to the legal field.

### v3

@@ 6.3 Analytical Reasoning
The traditional UI of legal databases merely lists groups of legal documents related to the currently displayed document. There are no means by which legal scholars can add evidence of their insights to that list. In VA, analytical reasoning is a knowledge-generation process that can have systematic aspects but is also serendipitous [shrinivasanSupportingAnalyticalReasoning2008]. Users must be aware of the evidence found throughout this process to increase the chance of insights. A prominent approach to capture such evidence is analytical provenance as suggested by [pirolli2005sensemaking] for their sense-making loop [pirolli2005sensemaking]. For that, [perez-messinaPersistentInteractionUserGenerated2024] suggest organizing analytical artifacts like annotations and their relationships [shrinivasanSupportingAnalyticalReasoning2008, perez-messinaPersistentInteractionUserGenerated2024].
In related text-based research practice, Overview allows investigative journalists to create tags that attach their findings to clusters [brehmerOverviewDesignAdoption2014]. [tianLitVisVisualAnalytics2023] add to this, allowing users to adapt clusters of literature interactively, adding comments about their understanding of topics [tianLitVisVisualAnalytics2023] (refer to 2, c). ForceSPIRE addresses visual text analytics more broadly, enabling experts to create text highlights that constitute a spatial layout of document relationships [endertSemanticInteractionVisual2012] (refer to 2, b).
Conveying the relationships within and between such clusters is central for analytical reasoning, so node-link diagrams are frequently employed as visualization methods. [lettieriLegalMacroscopeExperimenting2017] introduce the Norm Graph Navigator, allowing interactive exploration of legal relationships [lettieriLegalMacroscopeExperimenting2017]. More recently, [lacavaLawNetVizWebbasedSystem2022] integrate natural language querying into network navigation, enabling users to filter nodes efficiently [lacavaLawNetVizWebbasedSystem2022] (refer to 2, c). Since analytical artifacts externalize tacit knowledge as explicit [wangDefiningApplyingKnowledge2009], they can act as a countermeasure to the confirmation bias in hypothesis generation [pirolli2005sensemaking] and enable collaboration [perez-messinaPersistentInteractionUserGenerated2024], which can reinforce diverse opinions in legal decision-making. When persisted [perez-messinaPersistentInteractionUserGenerated2024], these artifacts can help users recall and replicate insights [raganCharacterizingProvenanceVisualization2016]. Analytical provenance additionally fosters transparency, which is particularly important in the legal domain, where the validity and reliability of insights must be meticulously documented. Integrating these aspects, for example, through 2D clustering for overview, graph networks for relations, and provenance trees for documentation, shows the significant potential of VA for jurisprudence.
@@ 7 Emergent Challenges for Visual Analytics Design
While numerous approaches exist for data navigation in law, knowledge representation and reasoning are often left to analysts to perform manually, as we have seen in the previous discussion. Crucially, no existing solution unifies all these areas of VA into a framework that exploits their synergetic effects. In the following, we explore the opportunities and challenges of developing such an integrated approach.
@@ 7.1 Addressing Jurisprudence
Figure 3: A mock-up of our proposed Visual Analytics workflow that illustrates the role of Concept Maps, Icicle Plots, and Treemaps in the different phases of the three-phase workflow. Cursors and arrows indicate possible user interactions.
@@
Traditionally, German legal scholars predominantly use keyword-based search interfaces, the prevalent UI paradigm [verlagc.h.beckohgBeckonline, jurisgmbhJuris]. However, this poses several challenges to legal scholars, hindering effective workflow support. While familiar, the paradigm limits scholars to primitive input, lacking the expressiveness required in the legal domain. Hence, we propose combining existing visualization techniques into a novel VA workflow (refer to 1) for jurisprudence with LLMs.
@@
As part of the initial workflow phase (refer to $W_{\texttt{P1}}$), the scholar scopes their research verbally. Instead of a syntactic keyword search, they articulate analysis goals using natural language (refer to 3). The system processes the verbalization and embeds the query into a semantic context by inferring references to legal entities such as laws, legal norms, and court rulings. We imagine an overview displaying clusters of legal concepts and individual documents relevant to the query, enabling scholars to leverage their domain expertise to refine the selection. For its implementation, the overview leverages a network visualization augmented with colored shapes that indicate clusters. For details, a scholar can preview individual documents on demand during selection, highlighting explanations for their relevance. After refining the scoping, legal scholars transition to the second workflow phase (refer to $W_{\texttt{P2}}$). In VA, the UI transitions into a detailed view upon selecting a cluster or document of interest (refer to 3). We propose juxtaposing the formal legal structure with a branching visualization such as an Icicle Plot. The former displays the legal documents of the scope with their hierarchy, enabling efficient navigation. Meanwhile, the Icicle Plot provides the contextualized path of the selected document in its hierarchy. Together, these visualizations address the legal domain’s unique requirement for simultaneous navigation across multiple hierarchies.
@@
During the analysis and interpretation, legal reasoning takes place, which must utilize the scholar’s domain knowledge. While the proposed VA techniques can initially visualize the explicit knowledge from the legal documents, they cannot draw on evolving domain knowledge. To achieve this, VA must provide interaction for users to articulate their tacit knowledge. Therefore, at any point during the workflow, if users find a relevant document missing from the current selection, they can establish a relationship to include it. The Treemap and Icicle Plot visualizations facilitate this interaction by integrating empty cells with suggestive icons as visual placeholders. By clicking these icons, the analyst can relate a selected document to others represented in the visualizations. These interactions reveal user preferences relevant to the subsequent recommendation of related legal documents. With these user preferences, we augment a knowledge graph (KG) of the legal documents selected during the first workflow phase. [yangLegalGNNLegalInformation2021] demonstrate the benefit of such an augmentation with LegalGNN, a framework for recommending legal documents based on a graph neural network trained with a legal KG and the contents of legal documents [yangLegalGNNLegalInformation2021]. We exploit the resulting recommendations to personalize the VA workflow with the expert’s domain knowledge. Further, we draw tacit knowledge from its externalization through analytical artifacts.
@@
Besides the Treemap and the Icicle Plot, we propose to add a concept map to the top side of the suggested user interface (refer to 3). The map allows users to synthesize and document evidence acquired throughout the legal reasoning process, forming a crucial part of the third workflow phase (refer to $W_{\texttt{P3}}$). Scholars capture and connect relevant artifacts within the concept map to organize their legal insights. Nodes represent primitive artifacts resembling individual legal entities, such as legal norms or court rulings. In contrast, compound artifacts describe the links that relate multiple primitives, representing causal connections in the legal realm. Based on these and the graph augmented during the previous workflow phase, we enhance documentation of the legal reasoning process through retrieval-augmentation generation (RAG). [edgeLocalGlobalGraph2025] propose GraphRAG, which performs RAG on the KG of text corpora [edgeLocalGlobalGraph2025]. The authors leverage the hierarchy of the KG to enable the summarization of documents at varying levels of abstraction. This approach aligns with the nature of legal corpora and can benefit the documentation of legal reasoning in jurisprudence. For instance, in analyzing racial profiling by the German Federal Police (see 5), a scholar might document that the police are responsible for securing train stations (§ 3 BPolG) and are permitted to identify individuals if necessary for fulfilling their duties (§ 23 BPolG). However, this authority must be balanced against the third-party effect of the German Basic Law (Art. 3 GG), which ensures equality and non-discrimination.
@@
Implementing the proposed VA workflow faces challenges. The suggested visualization techniques for VA may suffer from issues with visual scalability. For vast legal datasets, node-link diagrams, Treemaps, and Icicle Plots are of limited use. However, since the datasets are hierarchically structured, we can progressively unveil parts of the hierarchy to limit the information these visualization techniques need to convey. Further, leveraging progress in natural language understanding remains difficult: legal text data combines domain-specific language [ruethersRechtstheorieUndJuristische2022] and precise phraseology with intentional ambiguity at times [ruethersRechtstheorieUndJuristische2022]. This ambiguity allows for statutory interpretation [ruethersRechtstheorieUndJuristische2022], enabling laws to evolve alongside societal norms and values, but is challenging for LLMs. Although trained on vast datasets, LLMs still struggle to apply these nuances of legal reasoning, particularly open foundation models. For example, we found that Llama 3.1 70B still sometimes confuses § 212 StGB (Strafgesetzbuch, i.e., criminal code), which defines homicide for murder (§ 211 StGB). However, by definition of the criminal code, they are two strictly separate criminal offenses. While such models can describe the differences in great detail when explicitly asked, during reasoning, they can tend to collapse any distinctions without further warning. Equally, hallucinations, particularly during the subsumption phase, are critical as LLMs struggle with completeness. Therefore, it is crucial to perform extensive LLM fine-tuning to align with the specific language of the legal domain. This also presents ample opportunity for VA design to facilitate collaboration between human scholars and LLMs during the legal reasoning process.
@@ 7.2 Implications for Visual Analytics Design
The lingering challenges in data navigation, knowledge externalization, and analytical reasoning cannot be sufficiently addressed with AI and computational solutions alone. Therefore, we advocate for a human-is-the-loop design [endertHumanLoopNew2014], which positions the VA system as an integral part of the scholar’s workflow. This view contrasts designs that rely on the common human-in-the-loop [keimVisualAnalyticsDefinition2008, sachaKnowledgeGenerationModel2014] approach. While the latter assumes the domain expert to become a part of their VA system, the former prerequisites the VA system to become part of the domain expert and their workflow instead. To achieve this, the VA design process requires a user-centered approach that honors the workflows of the domain expert, supporting them throughout the sense-making loop [pirolli2005sensemaking, endertHumanLoopNew2014]. Therefore, we highlight the following key techniques and higher-level considerations regarding jurisprudence and generalization to other domains.
@@
User-Centered Design — The development of effective VA design requires adherence to user-centered design principles [gibbsBuildingBetterDigital2012, hinrichsSpeculativePracticesUtilizing2016]. These principles promote broader adoption of VA designs that meet specific audience needs. These needs vary across different application domains, valuing various visual qualities, such as page layout in law. While these aspects may not seem directly related to the effectiveness or efficiency of VA design, they can significantly influence data navigation in information spaces. This variation in user requirements presents ample opportunity for employing design studies or conducting pilot interviews with domain experts. These methodological approaches allow for gaining critical insights into the audience’s specific needs, workflows, and mental models, like the importance of tacit knowledge, the lack of contextualized search, and the three-phase research process, which we discussed in 4.
@@
Knowledge Externalization — In application domains, subject matter experts possess tacit knowledge regarding domain-specific problems. It encompasses implicit understanding, experiential insights, and intuitive reasoning processes not readily encoded in explicit knowledge. The externalization process captures these domain-specific analytical approaches in shareable analytical artifacts. If not articulated, VA design inherently lacks access to this knowledge, potentially compromising analytical effectiveness. While explicit knowledge may be available in substantial quantities through text documents, for example, integrating with tacit knowledge is imperative to enhance analytical outcomes [federicoRoleExplicitKnowledge2017]. VA design must strive to externalize tacit knowledge through user interaction, KGs, and analytical artifacts.
@@
Provenance Tracking — Analytical artifacts manifest knowledge externalization processes within VA systems. These artifacts serve as persistent representations of analytical reasoning and decision-making trajectories throughout the analytical workflow [perez-messinaPersistentInteractionUserGenerated2024]. Since users often struggle to maintain holistic awareness of past analytical steps, access to provenance information is essential for branching and revising the analysis process. VA design must facilitate the creation and storage of analytical artifacts through comprehensive provenance tracking, fostering the user’s awareness of the analysis process. Concurrently, these artifacts can be systematically utilized to guide subsequent analytical processes. VA design must exploit this recursive employment of analytical artifacts to create a self-reinforcing system that progressively refines the analytical workflow based on accumulated insights and knowledge.
@@
Natural Language Inference — While mining user interactions with VA interfaces can constitute guidance, it presents inherent methodological challenges in deducing higher-order analytical objectives [gotzCharacterizingUsersVisual2008]. Interaction patterns alone provide insufficient semantic context for accurately inferring analytical goals that guide user behavior. With recent advancements in LLMs, these models demonstrate sophisticated semantic parsing and intention recognition capabilities that can be leveraged to extract structured analytical objectives from natural language [zhaoLEVAUsingLarge2025]. VA design must enable users to articulate these analytical objectives through natural language interfaces to capture nuanced analytical intentions while going beyond a simple chat-based interface design.
@@
Unified Design — To achieve effectiveness and efficiency in addressing domain-specific problems, VA design must integrate the aforementioned key techniques. User-centered design facilitates identifying and formalizing domain-specific workflows, as demonstrated through our structured interviews with subject matter experts (see 4). Such approaches reveal critical insights into domain-specific analytical processes that may not be apparent through conventional design approaches. Likewise, knowledge externalization provides access to tacit knowledge, informing VA. Provenance tracking can effectively reflect domain-specific workflows by capturing the structure of analytical processes. Minimizing cognitive load by context switching and unifying analytical functionalities within a coherent interface allows one to optimize cognitive resource utilization and analytical performance.
@@ 7.3 Limitations
Our unified design approaches legal tech from the VA perspective, adopting a definition of visualization [cardReadingsInformationVisualization2007] from the visualization community in our literature analysis. In contrast, the legal community uses a broader definition [khalilQuoVadisVisuelle2014], leading to papers that visualize legal data but are not considered VA. While not our primary focus, we integrate such visualization concepts—familiar to legal scholars—into our VA design (see 6). For example, we employ Concept Maps, commonly used to represent cases and causal relations between norms [cantatoreMakingConnectionsIncorporating2016], to capture tacit knowledge (see 6.2) during legal reasoning. Thus, visualization concepts from law can enrich VA design and foster interdisciplinary research.
@@
Tasks — To enhance our understanding of how legal scholars operate, we describe different levels of tasks (see 3.2). These levels are supported by research on applications in legal tech, often emerging from the intersection of law and computer science. Consequently, the labels for these task levels resemble technical terms from VA, such as information retrieval. However, these tasks are grounded in the operations of scholars as described in literature from the legal domain (see 3.1). Additionally, these tasks are linked to the three-phase workflow (see 5), which is based on interviews with legal scholars (see 4).
@@
Expert Interviews — These results of the subject matter expert interviews (see 4) take common workflows and recurring challenges faced by German scholars from various legal disciplines into account. Still, they are not exhaustive, as further disciplines exist in the German legal system and beyond. To accommodate the differences between legal disciplines, we have decided on a semi-structured interview, where we adapt the follow-up questions to the interviewees’ backgrounds and experiences. Although we assume differences in the details, we expect the significant challenges distilled in 6 to be similar, as they concern all disciplines. We elicit our observations from subject matter experts who work exclusively with the German civil law system and European legislation. So far, we have not incorporated experiences from countries that practice civil law or other legal systems like common law (see 1).
Case Study — To illustrate the three-phase workflow design (see 5) and to critically discuss traditional UI for legal research in our case study (see 6), we have focused on the Bundespolizeigesetz (Federal Police Act). This specific legal research context highlights the challenges described in a broader context by participants during the interviews (see 4) using an example. Challenges arise from accessibility issues, navigational complexity, the inefficiency of the search interface, and the reliance on tacit knowledge, which extend to other laws and research contexts.
Generalization — This generalization allows our case study to transcend the specific example chosen and the German legal system. To illustrate the traditional UI used in jurisprudence as part of our case study (see 6), we use beck-online as an example. Although beck-online is just one of many commercial and scientific applications for jurisprudence, it virtually has a monopoly on legal research by practitioners. Additionally, it features a design similar to applications such as Westlaw [thomsonreutersWestlawClassic] and PACER [administrativeofficeoftheu.s.courtsPACERPublicAccess], which helps generalize our findings to case law. Further, our work lacks empirical, quantitative evaluations of the proposed VA workflows as a prototype, as empirical evaluations from other domains might not generalize to the legal field.